# DeCo-SGD: Joint Optimization of Delay Staleness and Gradient Compression Ratio for Distributed SGD

### v1

@@ 6 Related Work
Gradient Compression Strategy. Gradient compression methods are generally categorized into quantization alistarh2017qsgd ; bernstein2018signsgd ; seide20141 , sparsity aji2017sparse ; lin2017deep ; stich2018sparsified , hybrid approaches dryden2016communication ; jiang2018sketchml , and low-rank strategies vogels2019powersgd . Traditional schemes often rely on fixed compression parameters, which limit their adaptability to varying network and training conditions. To address this, recent works have introduced dynamic and adaptive strategies. DC2 abdelmoniem2021dc2 adjusts communication based on real-time network dynamics, while DAGC lu2023dagc assigns compression aggressiveness based on each worker’s data volume. However, neither of them considers end-to-end latency. FedCG jiang2023heterogeneity selects representative clients and adapts compression ratios according to their heterogeneity and capacity changes. L-GreCo markov2024greco formulates the per-layer compression allocation as a constrained optimization problem solved via dynamic programming to balance model size and error. However, neither of the above two methods takes into account the dynamic changes of the network environment.
Asynchronism. FedBuff nguyen2022federated uses a method of buffering asynchronous aggregation to improve communication efficiency and model performance. However, the buffering mechanism may increase memory overhead and delay the incorporation of fresh gradients. Mishchenko et al.mishchenko2022asynchronous rigorously proved theoretically and validated empirically that asynchronous SGD can achieve convergence comparable to mini-batch SGD, even in the presence of arbitrary gradient delays. CO2 sun2024co2 introduces local update and single-step asynchronous communication mechanisms to achieve complete overlap between communication and computing, but it assumes homogeneous device capabilities. To further address heterogeneity in node speed, Shadowheart SGD tyurinshadowheart dynamically adjusts the local update frequency and communication frequency of each node based on the different computing and communication speeds of each working node, thereby achieving the optimal time complexity in an asynchronous setting.
@@ 7 Conclusion
In this work, we address the challenge of efficient distributed training in high-latency, low-bandwidth and varying bandwidth networks, where D-SGD suffers significant throughput degradation. While gradient compression and delayed aggregation can mitigate bandwidth and latency issues, respectively, they introduce a complex trade-off among compression ratio, staleness, and model accuracy. To achieve the optimal balance during the whole training, we mathematically model it from two perspectives: time-to-iteration and iteration-to-accuracy. For the former, we derive the threshold compression ratio that maximizes computation-communication overlap under the given staleness and network conditions. For the latter, we introduce NVS to derive the convergence rate of DD-EF-SGD. We are the first to reveal that staleness exponentially amplifies the negative impact of gradient compression, filling the critical theoretical gap in DD-EF-SGD. Building on these insights, we propose DeCo-SGD, an adaptive optimizer that dynamically adjusts compression and staleness based on network conditions. Our design achieves up to 5.07×5.07$\times$5.07 × speed-up over D-SGD and 1.37×1.37$\times$1.37 × over the SOTA method. The code will be released upon publication.
@@ Appendix A Notation List
Table 2: Notation list.
Notation Description
$n$ number of workers
$\mathbf{T}$ total number of iterations
$f_{i}(\cdot)$ the local loss function of worker $i$
$\delta$ compression ratio ($0<\delta\leq 1$)
$\mathbf{C}_{\delta}(\cdot)$ the sparsification compressor with compression ratio $\delta$
$\mathbf{x}^{i}_{t}$ the local model parameter in the $t$-th iteration of the $i$-th worker
$\mathbf{x}_{t}$ the global model parameter in the $t$-th iteration
$\mathbf{g}^{i}_{t}$ the stochastic gradient of $\textbf{x}_{t}$ of the worker $i$
$\gamma$ the stepsize
$\tau$ the delayed staleness
$\mathbf{e}_{t}^{i}$ the local error term of the $i$-th worker at the $t$-th iteration
$L$ L-smoothness
$\mu$ $\mu$-strongly convexity if $f_{i}$ is strongly convex
$\zeta_{i}$ the data heterogenority of worker $i$
$\zeta$ the global data heterogenority, i.e., $\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}^{2}$
$\xi^{i}$ the stochastic gradient noise of worker $i$
$\sigma^{2}$ the upper variance bound of the stochastic gradient noise
$a$ the bandwidth (bits/s)
$b$ the network end-to-end latency (s)
$S_{g}$ the size of the gradient (bits)
$T_{comp}$ computation time per iteration (s)
$T_{avg}$ average end-to-end training time per iteration (s)
$T_{sum}$ total end-to-end training time (s)
$TS_{k}$ End time of the k-th computation(s)
$TM_{k}$ End time of the k-th transmission(s)
$TC_{k}$ End time of the k-th communication(s)
@@ B.1 Technical Results
$\lVert c_{1}+c_{2}\rVert^{2}\leq(1+\rho)\lVert a\rVert^{2}+(1+\rho^{-1})\lVert b\rVert^{2}.$ (12)
@@
$\langle c_{1},c_{2}\rangle^{2}\leq\rho\lVert a\rVert^{2}+\rho^{-1}\lVert b\rVert^{2}.$ (13)
@@
Lemma 3 (Lemma 27 of the work [19]).
@@
Lemma 4 (Lemma 25 of the work [19]).
@@
$r_{t+1}\leq(1-\min\{\gamma A,F\})r_{t}-B\gamma s_{t}+C\gamma^{2}+D\gamma^{3},$
for some positive constants $A,B>0$, $C,D\geq 0$, and for constant step-sizes $0<\gamma\leq\frac{1}{E}$, for $E\geq 0$, and for parameter $0<F\leq 1$. Then there exists a constant step-size $\gamma\leq\frac{1}{E}$ such that
$\frac{B}{W_{T}}\sum_{t=0}^{T}w_{t}s_{t}+\min\left\{A,\frac{F}{\gamma}\right\}r_{T+1}$ $\leq r_{0}\left(E+\frac{A}{F}\right)\exp\left[-\min\left\{\frac{A}{E},F\right\}(T+1)\right]$
$\quad+\frac{2C\ln\tau}{A(T+1)}+\frac{D\ln^{2}\tau}{A^{2}(T+1)^{2}}$
for $w_{t}:=(1-\min\{\gamma A,F\})^{-(t+1)}$, $W_{T}:=\sum_{t=0}^{T}w_{t}$ and
$\tau=\max\left\{\exp[1],\min\left\{\frac{A^{2}r_{0}(T+1)^{2}}{C},\frac{A^{3}r_{0}(T+1)^{3}}{D}\right\}\right\}$
@@
$\tilde{\mathcal{O}}\left(r_{0}\left(E+\frac{A}{F}\right)\exp\left[-\min\left\{\frac{A}{E},F\right\}T\right]+\frac{C}{AT}+\frac{D}{A^{2}T^{2}}\right),$
@@
$T=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{1}{F}\log\frac{1}{\epsilon}+\frac{E}{A}\log\frac{1}{\epsilon}\right)=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{1}{F}+\frac{E}{A}\right)$
@@ B.2 Proof of Theorem 1
$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\frac{\gamma}{2}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{t}\right)\right\|^{2}+\gamma L^{2}\mathbb{E}\left\|\tilde{B}_{t}\right\|^{2}+2\gamma L^{2}\mathbb{E}\left\|B_{t}\right\|^{2}+\frac{\gamma^{2}L\sigma^{2}}{2n}.$
@@
$\leq\mathbb{E}f\left(\tilde{\mathbf{x}}_{t}\right)-\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle+\frac{L}{2}\mathbb{E}\left\|\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right\|^{2}$
$\leq\mathbb{E}f\left(\tilde{\mathbf{x}}_{t}\right)-\underbrace{\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle}_{:=A_{1}}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\gamma^{2}\sigma^{2}}{2n}.$
@@
$\leq\frac{c_{1}\gamma L^{2}}{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\frac{\gamma}{2c_{1}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{c_{2}\gamma L^{2}}{2}\mathbb{E}\lVert B_{t}\rVert^{2}+\frac{\gamma}{2c_{2}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}-\gamma\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}.$
The last line is due to Jensen’s inequality (i.e., Eq. 13). We let $c_{1}=2$, $c_{2}=4$. By bringing Eq. LABEL:A-1 into Eq. 16, we complete the proof.
@@
Define $\{\tilde{B}_{t}\}_{t\geq 0}$ as in Eq. 3.2. Then with $\gamma\leq\frac{1}{4L\sqrt{\tau}}$, we have
@@
Proof. This lemma follows from the Lemma 20 in the previous work [14].
@@
Define $\{B_{t}\}_{t\geq 0}$ as in Eq. 3.2. Then by using the relative compressor, with $\gamma\leq\frac{1}{4L\sqrt{\phi/\delta}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, we have
$\sum_{t=0}^{T}\mathbb{E}\lVert B_{t}\rVert^{2}\leq\frac{1}{4L^{2}}\mathbb{E}\sum_{t=0}^{T}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi(T+1)(\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\frac{(1-\delta)(1+\rho)}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau}^{i}\rVert^{2}$
$+\gamma^{2}(1-\delta)((1+\rho^{-1})\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-\tau})\rVert^{2}+(1+\rho^{-1})\zeta^{2}+\sigma^{2}).$
This inequality is due to Lemma 2, Eq. 4, 3,12.
By unrolling the recurrence and let $\rho=\frac{2-\delta}{\delta}$, we can derive the upper bound:
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\sum_{t=0}^{T}[\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}]$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}\sum_{i=0}^{T}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{T}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
Choosing $\frac{4\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we have
@@
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. There exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\sqrt{\tau}},\frac{1}{4L\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}L}{n\epsilon^{2}}+\frac{L\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{L}{\epsilon}+\frac{L\sqrt{\tau}}{\epsilon}+\frac{\sqrt{\phi}L}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$
@@
$\frac{\gamma}{2}G_{T}\leq\mathbb{E}f(\hat{\mathbf{x}}_{0})-\mathbb{E}f(\hat{\mathbf{x}}_{T+1})+\gamma L^{2}(\sum_{i=0}^{T}\lVert\tilde{B}_{t}\rVert^{2}+\sum_{i=0}^{T}\lVert B_{t}\rVert^{2})+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}$
$\leq\mathbb{E}f(\mathbf{x}_{0})-f^{*}+\frac{\gamma}{4}G_{T}+\frac{\gamma}{8}G_{T}+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}+2\gamma^{3}L^{2}(T+1)(2\phi\frac{2}{\delta}\zeta^{2}+2\phi\sigma^{2}+\frac{2\tau\sigma^{2}}{n}).$
Divide both sides of the inequality by $4\gamma(1+T)$, we have:
$\frac{1}{8(1+T)}G_{T}\leq\frac{\mathbb{E}f(\mathbf{x}_{0})-f^{*}}{\gamma(T+1)}+\gamma\frac{L\sigma^{2}}{2n}+4\gamma^{2}L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n}).$
According to Lemma 3, we let $B=1/8$, $C=\frac{L\sigma^{2}}{2n}$, $D=L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n})$, $r_{0}=\mathbb{E}f(\mathbf{x}_{0})-f^{*}$. To ensure the right hand side of the above inequality is less than $\epsilon>0$, $T=\mathcal{O}(\frac{C}{\epsilon^{2}}+\frac{\sqrt{D}}{\epsilon^{3/2}}+\frac{E}{\epsilon})r_{0}$ iterations are sufficient, which completes the proof.

### v2

@@ 6 Related Work
Communication-efficient and adaptive training. To alleviate the communication bottlenecks in distributed training, gradient compression has become a standard practice. While early works focus on static compression strategies (e.g., fixed quantization bits or sparsity ratios), recent research has shifted towards dynamic and adaptive strategies enhance robustness against system heterogeneity or network dynamics. For instance, DAGC lu2023dagc assigns different compression ratios based on the local data volume of each worker to address communication problems in communication-constrained non-IID scenarios. FedCG Wu2021FedCGLC selects representative clients and adapts compression ratios according to their heterogeneity and capacity changes to make the system robust to device heterogeneity. HCEF HCEF theoretically analyzes the impact of local update frequency and gradient compression on the convergence bound in Cooperative Federated Edge Learning. It develops an online algorithm to dynamically determine these two parameters to minimize training latency and energy communication. HGC hu2024practical compresses both uplink and downlink communications in federated learning, removing redundancy in the training process to greatly reduce data transfer while keeping the model performance. DC2 dc2 adaptively controls the gradient compression ratio based on network delays, effectively managing network variations to strike a better balance between accelerating distributed training and preserving the model accuracy. These methods, however, are largely designed for bandwidth efficiency and remain insensitive to the fixed latency. This limitation leaves model performance vulnerable in challenging network environments where high latency coexists with limited bandwidth.
Latency-tolerant training. To enhance the tolerance of training to high latency, relaxation of the strict synchronization barrier is a common approach. Fully asynchronous methods are primarily designed to address the straggler effect caused by heterogeneous device capabilities. For example, FedAC zhang2024fedac incorporates look-ahead momentum to correct updates in buffered aggregation, while FedFa xu2024fedfa utilizes an update buffer to enable fully asynchronous updates. To address computational heterogeneity, Shadowheart SGD tyurinshadowheart dynamically adjusts local update and communication frequencies, and P2P-AFL sad2025towards proposes a decentralized model fusion mechanism. While these approaches effectively reduce waiting times, However, the fully asynchronous strategy may introduce extremely high staleness, leading to model accuracy degradation. It is also unsuitable for WAN-based training, which has relatively fixed latency. To address this, DGA DGA introduces delayed aggregation into FL, attempting to solve the fixed end-to-end latency issue in FL. However, it does not account for communication bottlenecks. CO2 sun2024co2 introduces local update and delayed aggregation with $\tau{=}1$ to achieve complete overlap between communication and computing. But it focuses on high-speed interconnects and is not suitable for bad network conditions, which require adaptive adjustment of the staleness to overlap communication and computation.
Distributed training in WANs. WAN-based training faces the challenges of high latency and low, varying bandwidth. Early efforts, such as MMOptree liu2022accelerating and NETSTORM li2024accelerating, optimize the logical topology to reduce hop counts. However, topology optimization alone cannot fully eliminate the communication bottlenecks and high end-to-end latency. StellaTrain lim2024accelerating integrates cache-aware compression and sparse optimization for geo-distributed clusters. However, StellaTrain lacks rigorous theoretical guidance regarding the trade-off between compression rate and staleness, relying instead on empirical tuning. Consequently, it may fail to maintain optimal convergence rates under the extreme and fluctuating network conditions typical of WANs.
@@ 7 Conclusion
In this work, we address the challenge of efficient distributed training in high-latency, low-bandwidth and varying bandwidth networks, where D-SGD suffers significant throughput degradation. While gradient compression and delayed aggregation can mitigate bandwidth and latency issues, respectively, they introduce a complex trade-off among compression ratio, staleness, and model accuracy. To achieve the optimal balance during the whole training, we mathematically model it from two perspectives: time-to-iteration and iteration-to-accuracy. For the former, we derive the threshold compression ratio that maximizes computation-communication overlap under the given staleness and network conditions. For the latter, we introduce NVS to derive the convergence rate of DD-EF-SGD. We are the first to reveal that staleness exponentially amplifies the negative impact of gradient compression, filling the critical theoretical gap in DD-EF-SGD. Building on these insights, we propose DeCo-SGD, an adaptive optimizer that dynamically adjusts compression and staleness based on network conditions. Our design achieves up to $5.07\times$ speed-up over D-SGD and $1.37\times$ over the SOTA method. The code will be released upon publication.
@@ B.1 Technical Results
$\lVert c_{1}+c_{2}\rVert^{2}\leq(1+\rho)\lVert c_{1}\rVert^{2}+(1+\rho^{-1})\lVert c_{2}\rVert^{2}.$ (12)
@@
$\langle c_{1},c_{2}\rangle^{2}\leq\rho\lVert c_{1}\rVert^{2}+\rho^{-1}\lVert c_{2}\rVert^{2}.$ (13)
@@
Lemma 3 (Lemma 27 of the work [stich2020communication]).
@@
Lemma 4 (Lemma 25 of the work [stich2020communication]).
@@
$r_{t+1}\leq(1-\gamma A)r_{t}-B\gamma s_{t}+C\gamma^{2}+D\gamma^{3},$
for some positive constants $A,B>0$, $C,D\geq 0$, and for constant step-sizes $0<\gamma\leq\frac{1}{E}$, for $E\geq 0$. Then we have
$\frac{B}{W_{T}}\sum_{t=0}^{T}w_{t}s_{t}+Ar_{T+1}\leq r_{0}E\exp\left[-\frac{A}{E}(T+1)\right]+\frac{2C\ln\tau}{A(T+1)}+\frac{D\ln^{2}\tau}{A^{2}(T+1)^{2}},$
for $w_{t}:=(1-\gamma A)^{-(t+1)}$, $W_{T}:=\sum_{t=0}^{T}w_{t}$ and $\tau=\max\left\{\exp[1],\min\left\{\frac{A^{2}r_{0}(T+1)^{2}}{C},\frac{A^{3}r_{0}(T+1)^{3}}{D}\right\}\right\}$.
@@
$\tilde{\mathcal{O}}\left(r_{0}E\exp\left[-\frac{A}{E}\cdot T\right]+\frac{C}{AT}+\frac{D}{A^{2}T^{2}}\right),$
@@
$T=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{E}{A}\log\frac{1}{\epsilon}\right)=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{E}{A}\right)$
@@ B.2 Proof of Theorem 1
$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\frac{\gamma}{2}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{t}\right)\right\|^{2}+\gamma L^{2}(2\mathbb{E}\left\|\tilde{B}_{t}\right\|^{2}+\mathbb{E}\left\|B_{t}\right\|^{2})+\frac{\gamma^{2}L\sigma^{2}}{2n}.$
@@
$\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle+\frac{L}{2}\mathbb{E}\left\|\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right\|^{2}$
$\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\underbrace{\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle}_{:=A_{1}}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\gamma^{2}\sigma^{2}}{2n}.$
@@
$\leq\frac{c_{1}\gamma L^{2}}{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\frac{\gamma}{2c_{1}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{c_{2}\gamma L^{2}}{2}\mathbb{E}\lVert B_{t}\rVert^{2}+\frac{\gamma}{2c_{2}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}-\gamma\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}$
$\leq-\frac{5\gamma}{8}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\gamma L^{2}(2\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
The third line is due to Jensen’s inequality (i.e., Eq. 13). In the fourth line, we let $c_{1}=4$, $c_{2}=2$. Bringing Eq. LABEL:A-1 into Eq. 16, we complete the proof.
@@
Define $\{\tilde{B}_{t}\}_{t\geq 0}$ as in Eq. 6. Then with $\gamma\leq\frac{1}{4L\tau}$, we have
@@
Proof.
$\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$ $=\gamma^{2}\mathbb{E}\lVert\sum_{j=1}^{\tau}(\nabla f(\mathbf{x}_{t-j})+\frac{1}{n}\sum_{i=1}^{n}\xi^{i})\rVert^{2}$
$\leq\gamma^{2}(\tau\sum_{j=1}^{\tau}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-j})\rVert^{2}+\frac{1}{n^{2}}\cdot n\tau\sigma^{2}).$
The inequality is due to Jensen’s inequality and Eq. 2. Then we have
$\sum_{t=0}^{T}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$ $\leq\gamma^{2}\tau\sum_{t=0}^{T}\sum_{j=1}^{\tau}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-j})\rVert^{2}+\frac{\gamma^{2}\tau\sigma^{2}(1+T)}{n}$ (18)
$\leq\gamma^{2}\tau\cdot\tau\sum_{t=0}^{T-\tau}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{\gamma^{2}\tau\sigma^{2}(1+T)}{n}$
$\leq\gamma^{2}\tau^{2}\sum_{t=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{\gamma^{2}\tau\sigma^{2}(1+T)}{n}.$
The second line is due to the fact that the term $\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}$ appears at most $\tau$ times. Let $\gamma\leq\frac{1}{4L\tau}$, we complete the proof.
@@
Define $\{B_{t}\}_{t\geq 0}$ as in Eq. 6. Then by using the relative compressor, with $\gamma\leq\frac{1}{4LZ\sqrt{\phi/\delta}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, we have
$\sum_{t=0}^{T}\mathbb{E}\lVert B_{t}\rVert^{2}\leq\frac{1}{4L^{2}}\sum_{t=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi(T+1)(\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\frac{(1-\delta)(1+\rho)}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau}^{i}\rVert^{2}$ (19)
$\quad+\gamma^{2}(1-\delta)((1+\rho^{-1})Z^{2}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-\tau})\rVert^{2}+(1+\rho^{-1})\zeta^{2}+\sigma^{2})$
$\leq\frac{1}{n}(1-\frac{\delta}{2})\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau}^{i}\rVert^{2}+\gamma^{2}(1-\delta)(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-\tau})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
The first inequality is due to Lemma 2, Eq. 3, 2 and 12. In the second line, we let $\rho=\frac{\delta}{2-\delta}$, having $1+\rho^{-1}=\frac{2}{\delta}$ and $(1-\delta)(1+\rho)=1-\frac{\delta}{2-\delta}<1-\frac{\delta}{2}$.
By unrolling the recurrence, we can derive the upper bound:
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}(1-\frac{\delta}{2})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}),$
$\sum_{t=0}^{T}[\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}]$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}\sum_{i=0}^{T}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{T}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
Choosing $\frac{4\gamma^{2}(1-\delta)Z^{2}}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we have
@@
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. There exists a stepsize $\gamma\leq\min\{\frac{1}{4L\tau},\frac{1}{4LZ\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon}+\frac{Z\sqrt{\phi}}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$
@@
$\frac{\gamma}{2}G_{T}$ $\leq\mathbb{E}f(\hat{\mathbf{x}}_{0})-\mathbb{E}f(\hat{\mathbf{x}}_{T+1})+\gamma L^{2}(2\sum_{i=0}^{T}\lVert\tilde{B}_{t}\rVert^{2}+\sum_{i=0}^{T}\lVert B_{t}\rVert^{2})+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}$
$\leq\mathbb{E}f(\mathbf{x}_{0})-f^{*}+\frac{\gamma}{4}G_{T}+\frac{\gamma}{8}G_{T}+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}+\gamma^{3}L^{2}(T+1)(2\phi\frac{2}{\delta}\zeta^{2}+2\phi\sigma^{2}+\frac{2\tau\sigma^{2}}{n}).$
Divide both sides of the inequality by $4\gamma(1+T)$, and we have:
$\frac{1}{8(1+T)}G_{T}\leq\frac{\mathbb{E}f(\mathbf{x}_{0})-f^{*}}{\gamma(T+1)}+\gamma\frac{L\sigma^{2}}{2n}+2\gamma^{2}L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n}).$
We do not believe that DD-EF-SGDwill degrade into D-SGD, meaning that $\tau=0$ and $\delta=1$ cannot coexist at the same time. In this way, we have $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\tau},\frac{1}{4LZ\sqrt{\phi/\delta}}\}\leq\min\{\frac{1}{4L\tau},\frac{1}{4LZ\sqrt{\phi/\delta}}\}$.
According to Lemma 3, we let $B=1/8$, $C=\frac{L\sigma^{2}}{2n}$, $D=L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n})$, $E=\max\{4L\tau,4LZ\sqrt{\phi/\delta}\}$, $r_{0}=\mathbb{E}f(\mathbf{x}_{0})-f^{*}$. To ensure the right hand side of the above inequality is less than $\epsilon>0$, $T=\mathcal{O}(\frac{C}{\epsilon^{2}}+\frac{\sqrt{D}}{\epsilon^{3/2}}+\frac{E}{\epsilon})\cdot r_{0}$ iterations are sufficient, which completes the proof.